A body moving with an initial velocity 6ms~-¹ accelerates uniformly at a rate of 50ms~-¹.What is the distance covered during this period

1
Expert's answer
2022-03-14T09:45:53-0400

The distance traveled by a body

"d=v_0t+\frac{at^2}{2}"

For the time "t=4\:\rm s" we obtain

"d=6*4+\frac{50*4^2}{2}=124\:\rm m"